Question 5 of 10 (1 point) | Attempt 1 of Unlimited | View question in a popup 11.2 Section Exercise 14 (p-value, table) HurrPart 2 of 5 Compute the value of the test statistic. Round the answer to two decimal places.
Answer #1

The pooled proportion here is computed as:

P = (76 + 24) / (360 + 180) = 100/540 = 0.1852

The standard error here is computed as

SE = \sqrt{P(1-P)(\frac{1}{n_1} + \frac{1}{n_2})} = \sqrt{0.1852(1 - 0.1852)(\frac{1}{360} + \frac{1}{180})}

SE = 0.0355

The sample proportions now are computed here as:
p1 = 76/360 = 0.2111
p2 = 24/180 = 0.1333

Therefore the test statistic here is computed as:

z^* = \frac{p_1 - p_2}{SE} = \frac{0.2111 - 0.1333}{0.0355} = 2.19

Therefore 2.19 is the test statistic value here.

As this is a one tailed test, the p-value here is obtained from the standard normal tables as:
p = P(Z > 2.19) = 0.0143

As the p-value here is 0.0143 < 0.1 which is the level of significance here, therefore the test is significant here and we reject the null hypothesis here. Therefore we have sufficient evidence here that the first population proportion is more than the second.